OpenStudy (anonymous):

plz answer for year 7 student:( question: 16+-25

OpenStudy (anonymous):

this? 16 + (-25) = 16 - 25 = -9

OpenStudy (anonymous):

thank you so much ur so nice

OpenStudy (anonymous):

anyone else can plz explain more

OpenStudy (anonymous):

let me elaborate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

by the way im in yr 7 so if u could explain in that form

OpenStudy (anonymous):

i'm not sure what way you want me to explain but i'll keep it simple. when you have a (+) and a (+) sign together, you get a (+)number. e.g. 4+4=8

OpenStudy (anonymous):

ok.. im getting it

OpenStudy (anonymous):

when you have two negative numbers, the answer turns into a positive. always. this is just a rule of math. e.g. (-4)(-2)=+8 or (-6)/(-3)=+2

OpenStudy (anonymous):

make sense so far?

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I get it

OpenStudy (anonymous):

better

OpenStudy (anonymous):

so when there is a (-) and a (+) number, the answer is always (-). e.g. (5)(-4)=-20 or (-8)/(2)=-4

OpenStudy (anonymous):

oooh now I get it

OpenStudy (anonymous):

thx so much

OpenStudy (anonymous):

no problem